# To Rely or Not to Rely? Evaluating Interventions for Appropriate Reliance on Large Language Models

### v1

@@ Abstract.
As Large Language Models become integral to decision-making, optimism about their power is tempered with concern over their errors. Users may over-rely on LLM advice that is confidently stated but wrong, or under-rely due to mistrust. Reliance interventions have been developed to help users of LLMs, but they lack rigorous evaluation for appropriate reliance. We benchmark the performance of three relevant interventions by conducting a randomized online experiment with 400 participants attempting two challenging tasks: LSAT logical reasoning and image-based numerical estimation. For each question, participants first answered independently, then received LLM advice modified by one of three reliance interventions and answered the question again. Our findings indicate that while interventions reduce over-reliance, they generally fail to improve appropriate reliance. Furthermore, people became more confident after making incorrect reliance decisions in certain contexts, demonstrating poor calibration. Based on our findings, we discuss implications for designing effective reliance interventions in human-LLM collaboration.
@@ 1. Introduction
As people increasingly incorporate Large Language Models (LLMs) into their decision-making workflows, there is cause for both tremendous optimism and skeptical concern (Zhao et al., 2023; Eigner and Händler, 2024). LLMs encode remarkably vast amounts of knowledge; can communicate with users in clear, natural language; and have the potential to improve both the quality and efficiency of users’ decision-making (Tankelevitch et al., 2024). However, people can over-rely on LLM advice (Spatharioti et al., 2023; Kabir et al., 2024; Si et al., 2023) or, in contrast, under-rely on it if they mistrust their LLM partner (Zhang, 2023). LLMs are prone to hallucination and error, yet often present their outputs with impressive detail and confidence (Zhang et al., 2023; Chen and Shu, 2023). In contrast, algorithm aversion—the tendency to distrust algorithmic agents—can lead users to ignore helpful recommendations due to general distrust towards the technology (Dietvorst et al., 2015; Burton et al., 2020).
@@
Conscious effort must therefore be placed on helping users calibrate appropriate levels of trust and reliance on LLM advice (Zhang et al., 2020; Zhou et al., 2024). Neither over-reliance nor under-reliance is ideal; instead, the focus should be to achieve appropriate reliance (Schemmer et al., 2022; Ma et al., 2023; He et al., 2023). Research in the tangential field of human-AI collaboration has explored methods like explanations and calibration techniques to improve joint decision-making by enhancing AI transparency and verifiability, though results often show mixed outcomes and trade-offs (Vasconcelos et al., 2023b; Liao and Varshney, 2021; Miller, 2019; Schoeffer et al., 2024; Bansal et al., 2021). In this work, we adapt rigorous measures of appropriate reliance from prior works in human-AI collaboration (Schemmer et al., 2023) to benchmark the impact of several LLM reliance interventions in human-LLM collaboration.
@@
Due to the risk of hallucinations, reliance and calibration interventions in LLMs are often motivated by curbing over-relying on wrongful information (Kim et al., 2024b; Si et al., 2023; Spatharioti et al., 2023), however recent attention has also shifted to optimizing for appropriate reliance (Zhou et al., 2024; Zhang, 2023; Collins et al., 2024; Do et al., 2024). Reliance interventions that involve communicating the capabilities and confidence of an LLM to its users (Spatharioti et al., 2023), or introducing various types of frictions that the user must bypass to receive the advice (Collins et al., 2024), are being actively developed and evaluated. Since interventions are evolving rapidly, many of these methods still lack comprehensive evaluations of appropriate reliance in real users, especially those focusing on technical contributions (Lin et al., 2022; Becker and Soatto, 2024; Bowman et al., 2022). Furthermore, people’s subjective perceptions—such as their self-confidence—can have a complex relationship with their reliance behaviour (Kim et al., 2024b; Cao and Huang, 2022; Lu and Yin, 2021; Si et al., 2023), which warrants further exploration. Finally, the impact of different intervention types on reliance and perceptions remains unclear, as existing studies do not benchmark multiple interventions. We seek to address these gaps by asking the following research questions:
@@ 2.1. Appropriate Reliance on AI
Factors that influence reliance include the expertise of the user in the task, their cognitive engagement in the task, and the information presented by the AI (Chen et al., 2023; He et al., 2023; Gajos and Mamykina, 2022). On the individual level, people may also be averse or appreciative towards algorithms, perhaps through pre-conceptions or judgments formed after seeing the AI’s performance (Kim et al., 2023; Dietvorst et al., 2015; Lu et al., 2023; Nourani et al., 2021; Ma et al., 2023). A number of strategies have been developed to augment users’ engagement and to provide them with more information of the AI decision, such as through providing explanations (Reingold et al., 2024; Vasconcelos et al., 2023b; Zhang et al., 2020; Schoeffer et al., 2024). The best interventions for appropriate reliance should be designed to support the users’ abilities to verify the AI’s recommendation, not necessarily to provide a fully transparent system (Fok and Weld, 2023; Poursabzi-Sangdeh et al., 2021; Wang and Yin, 2021). We adapt appropriate reliance metrics (Schemmer et al., 2023) to assess the ability of LLM interventions to help users verify outputs and increase appropriate reliance.
@@ 2.2. LLMs as Decision Aids
Recent research has investigated the role of LLMs in human-LLM collaborations for applications such as writing (Liang et al., 2024), communication (Fu et al., 2024), coding (Kazemitabaar et al., 2023; Nam et al., 2024), education (Lo, 2023; Kumar et al., 2023b; Bo et al., 2024), fact-checking (Wolfe and Mitra, 2024; Kim et al., 2024a; Si et al., 2023), information search (Spatharioti et al., 2023; Salikutluk et al., 2023; Sharma et al., 2024), and other specialized domains (Choudhury and Shamszare, 2023; Cheong et al., 2024; Yang et al., 2023). In these tasks, the LLM can provide guidance and information for the user to assist in their decision-making, learning, and creations (Zhang, 2023). They can also offer antagonistic advice to induce the user into considering alternative options (Chiang et al., 2024; Ma et al., 2024a; Cai et al., 2024). Even when the user is just reading without explicitly searching for advice, the opinions embedded in the LLM outputs can be influence the user’s own opinions (Ferguson et al., 2023; Sharma et al., 2024; Krügel et al., 2023). In technical AI research, the quality of LLM responses has been improved through methods such as chain-of-thought with intermediate reasoning steps (Wei et al., 2022), debating amongst multiple LLMs (Khan et al., 2024; Kim et al., 2024a), and enforcing self-consistency in outputs (Wang et al., 2022; Becker and Soatto, 2024).
In many cases, the human-LLM team can reach better performance as a team than individually, tapping into the strength of their combined knowledge and skill sets (Bowman et al., 2022). However, the question of how to induce appropriate reliance is still open. While using an LLM for information search can be just as efficient as a search engine, the user’s ability to distinguish incorrect information is diminished, resulting in a higher rate of over-reliance on incorrect information (Si et al., 2023; Xu et al., 2023). On the other hand, aversion to LLM advice has also been observed, resulting in users losing performance gains that they could have achieved through relying on the LLM (Zhang, 2023).
@@ 3. Methodology
We designed a large-scale, randomized, and incentivized experiment to evaluate the efficacy of three reliance interventions applied to the outputs of GPT-4o111We selected GPT-4o (https://openai.com/index/hello-gpt-4o/) as it is the most state-of-art LLM at the time of experimentation. in two question-answering tasks. To answer our research questions, we will compare the effect of the interventions on performance, reliance behaviour, confidence calibration, and time of engagement against the baseline LLM advice in the Control condition. In this section, we outline our choices of tasks, procedure, measures, and conditions.
@@ 3.1. Question-Answering Tasks
LSAT Logical Reasoning. Based on preliminary results from piloting multiple-choice questions from several domains, we found that logical reasoning questions best balanced difficulty with accessibility. We selected LSAT logical reasoning questions as the domain for the multiple-choice task and adapted two multiple-choice LSAT logical reasoning questions, both labeled ‘difficult’, for the first task. Each LSAT question has five options, of which exactly one is correct. We generated both correct and incorrect LLM advice for each question for use during the experiment. The questions and some sample LLM advice can be found in Figures A.1-A.2 in Appendix A. Correct answers were determined using LSAT answer keys.
@@
While LSAT questions are highly detailed, they are designed to test for aptitude in law school rather than knowledge of the legal domain. As such, they do not contain undefined jargon, nor do they require domain knowledge (satisfying D1). In addition to reasoning, LSAT Logical Reasoning questions also test analytical and reading comprehension skills, making them non-trivial to answer (satisfying D2). Furthermore, we reworded a question so that entering it into a search engine would not result in the correct answer (satisfying D3). Finally, they lend our study further external validity by emulating situations where people use LLMs to assist them with text comprehension or reasoning, in addition to people studying for tests such as the LSAT. While the MMLU dataset is another popular multiple-choice question benchmark for LLMs (Hendrycks et al., 2020), we did not find it suitable since it is skewed towards fact-based, know-it-or-you-don’t questions.
@@
Numerical Estimation. While most previous studies in human-AI collaboration focus on multiple-choice or binary outcome questions, we believe that introducing a numerical estimation task adds a valuable new dimension to this research space. Among the various forms of numerical estimation, such as predicting values in regression tasks, we opted for an image-based format based on a classic carnival game: estimating the quantity of beans in a glass jar. We posit that this is a gamified task that is familiar to most people and does not require advanced mathematics (satisfying D1). Despite the simplicity of the estimation process, it is still challenging to achieve accurate results (satisfying D2). Lastly, since we took the pictures and many of the beans are occluded, it is impossible to use external assistance to “cheat” on this task (satisfying D3).
@@
For the experiment, we captured photos of glass jars with manually-counted quantities of beans from both side perspectives and bottom perspectives. One set featured a smaller jar containing 403 chickpeas, while the other contained a larger jar with 856 jellybeans. The pictures and sample LLM advice can be found in Figures A.3-A.4 in Appendix A. Unlike in multiple-choice tasks, where advice can be objectively classified as good or bad, the goodness of the advice in the numerical estimation setting depends on the participant’s initial estimate. For example, while advice recommending a number that is far lower than the ground truth is objectively wrong, it is still better than an even lower participant estimate. To account for this variability, we provided five distinct levels of LLM advice relative to the ground truth quantity: very low, low, good, high, and very high. We chose these levels of advice based on results from initial pilot experiments to balance the proportion of bad advice with good advice. Visualizations of the advice levels with the distribution of participant guesses are shown in Figure C.7 in Appendix C.
@@ 3.2. Procedure
In the experiment, each participant was randomly assigned to one of four conditions and then answered two questions from the LSAT task and two questions from the Numerical Estimation task, randomized so that the LSAT phase and Numerical Estimation phase were equally likely to appear first. The orders of questions within each category were also randomized for each participant. Participants received pre-generated advice from GPT-4o. This advice was randomly sampled for each question, meaning participants could receive any mix of good or bad advice. After completing the tasks, participants completed a short post-survey. See Figure 2(a) for the full diagram of the experimental flow and Figure 2(b) for sample advice from all four conditions.
@@
Incentivization: Real-world decision-making involves balancing risk and reward. To simulate these stakes and induce incentive-compatible behaviour, we introduced a bonus, amounting to a maximum of 20% of the total participation pay. Participants earn rewards for correct answers, with two opportunities for bonus pay: one before, and one after receiving LLM-generated advice. In the LSAT questions, participants received a bonus if they selected the correct answer, while for the numerical estimation task, participants earned a bonus if they guessed within 10% of the true count.
@@ 3.3. Measures
What does it mean to rely on LLM advice? In multiple-choice questions, we follow previous work by defining reliance as changing one’s answer to the LLM’s suggestion. In estimation questions, we define reliance as an answer with positive weight-of-advice (WoA), where WoA is calculated as . We further define the reliance outcomes as falling into one of four categories:
@@
Negative LLM Reliance (NL): the participant changes their original answer to follow the LLM’s advice, but worsens or fails to improve their performance.
Negative Self-Reliance (NS): the participant does not follow the LLM’s advice, but doing so would have improved their performance.
PL and PS can be grouped together as Correct Reliance Behaviours, while NL and NS are Wrong Reliance Behaviours. NL represents over-reliance on LLM advice (the participant took the advice when it worsens or fails to improve improve their performance) and NS represents under-reliance (the participant didn’t take the advice but should have). We note that PL is possible even if the LLM advice is incorrect, as long as there is a notion of relative error—such as switching to a less wrong answer suggested by the LLM, which applies to the Numerical Estimation task and can extend to other real-world tasks. In contrast, the LSAT task lacks relative scoring across options, treating all incorrect answers as equally wrong; so relying on incorrect advice counts as over-reliance. This disambiguation is made since the LLM advice would not improve performance, so it would be preferable for the user to engage in the task and self-rely, allowing them to learn from the experience (Gajos and Mamykina, 2022). From these behaviour classifications, we compute the following appropriate reliance metrics:
Relative LLM Reliance (RLR) = Rate of appropriate reliance on correct LLM advice.
Relative Self-Reliance (RSR) = Rate of appropriate self-reliance when the LLM advice is wrong.
Appropriate Reliance Ratio (ARR) = Ratio of appropriate reliance on correct advice to over-reliance on wrong advice, where:
ARR ¡ 1: Higher rate of over-reliance than positive reliance.
@@
ARR ¿ 1: Higher rate of positive reliance than over-reliance.
@@
Confidence and Subjective Perceptions. For each question, we collect the participant’s confidence in the answer they gave prior to receiving LLM advice (Initial Confidence), and their confidence in the answer they gave after (Final Confidence) receiving the LLM-generated advice. Confidence is measured using a unipolar Likert scale from 1 (highly uncertain) to 5 (highly certain), which allows us to calculate the Change in Confidence from the first to second stage. Well-calibrated confidence should result in lower confidence scores when participants make incorrect reliance decisions. We compare the change in confidence between correct and wrong reliance outcomes for each of the conditions, where statistical differences are tested with Mann-Whitney U-test for non-normal distributions with significance threshold .
@@
ChatGPT’s advice made me more confident.222This subjective confidence measure is separate from the question-level confidence measure (the actual target of RQ2).
@@
Task Performance. Since our experiment is not designed to emulate the realistic performance of LLMs in tasks, we do not particularly emphasize the human-LLM collaborative task performance. However, we explore the initial error from participants independently answering questions, and compare this with their updated error after re-answering questions with LLM advice. We also compute a Best Possible Performance metric, which simulates the outcomes as if the human made the correct reliance decision in every single question, representing the best possible accuracy of a human-LLM team.
Time of Engagement. To understand participant engagement with the LLM’s advice, we measure the amount of time that participants spend on answering each stage of question-answering. From these two time measurements, we calculate the relative time as the ratio between time taken to engage with the advice, and the time taken to answer the initial question (Relative Advice Time).

### v2

@@ Abstract.
As Large Language Models become integral to decision-making, optimism about their power is tempered with concern over their errors. Users may over-rely on LLM advice that is confidently stated but wrong, or under-rely due to mistrust. Reliance interventions have been developed to help users of LLMs, but they lack rigorous evaluation for appropriate reliance. We benchmark the performance of three relevant interventions by conducting a randomized online experiment with 400 participants attempting two challenging tasks: LSAT logical reasoning and image-based numerical estimation. For each question, participants first answered independently, then received LLM advice modified by one of three reliance interventions and answered the question again. Our findings indicate that while interventions reduce over-reliance, they generally fail to improve appropriate reliance. Furthermore, people became more confident after making wrong reliance decisions in certain contexts, demonstrating poor calibration. Based on our findings, we discuss implications for designing effective reliance interventions in human-LLM collaboration.
Visual representation of four conditions. Each condition is represented by a text bubble coming from the right, with a text box on the left side. From left to right, the four conditions are Baseline, Reliance Disclaimer, Uncertainty Highlighting, and Implicit Answer. The textbox for Baseline contains “The answer is A” followed by lines representing unspecified text. The textbox for Reliance Disclaimer contains “The answer is A” followed by lines representing unspecified text, followed by “Remember to verify the information!” The textbox for Uncertainty Highlighting contains “The answer is A” followed by lines representing unspecified text, parts of which are highlighted in light pink or dark pink. The textbox for Baseline contains lines representing unspecified text.
@@ 1. Introduction
As people increasingly incorporate Large Language Models (LLMs) into their decision-making workflows, there is cause for both tremendous optimism and skeptical concern (Zhao et al., 2023; Eigner and Händler, 2024). LLMs encode remarkably vast amounts of knowledge; can communicate with users in clear, natural language; and have the potential to improve both the quality and efficiency of users’ decision-making (Tankelevitch et al., 2024). While LLMs are prone to hallucination and error, they can still mislead users into a false sense of trust with their impressive detail and confidence (Zhang et al., 2023; Chen and Shu, 2023). On the other hand, algorithm aversion can cause users to also ignore helpful recommendations due to a general distrust towards technology (Dietvorst et al., 2015; Burton et al., 2020). Consequently, both over-reliance (Spatharioti et al., 2023; Kabir et al., 2024; Si et al., 2023) and under-reliance present significant yet contrasting challenges in human-LLM collaboration (Zhang, 2023).
@@
Conscious effort must be placed on helping users calibrate appropriate levels of trust and reliance on LLM advice (Zhang et al., 2020; Zhou et al., 2024). Neither over-reliance nor under-reliance is ideal; instead, the focus should be to achieve appropriate reliance (Schemmer et al., 2022; Ma et al., 2023; He et al., 2023). Due to the risk of hallucinations, reliance interventions in LLMs are often motivated by curbing over-relying on wrongful information (Kim et al., 2024b; Si et al., 2023; Spatharioti et al., 2023), however recent attention has also shifted to optimizing for appropriate reliance as well (Zhou et al., 2024; Zhang, 2023; Collins et al., 2024; Do et al., 2024). In this work, we adapt rigorous measures of appropriate reliance from prior works in human-AI collaboration (Schemmer et al., 2023) to benchmark the impact of several LLM reliance interventions in question-answering tasks.
@@
The tangential field of human-AI collaboration has explored methods like explanations and calibration techniques to improve joint decision-making by enhancing AI transparency and verifiability, though results often show mixed outcomes and trade-offs (Vasconcelos et al., 2023b; Liao and Varshney, 2021; Miller, 2019; Schoeffer et al., 2024; Bansal et al., 2021). Reliance and calibration interventions for LLMs can involve communicating the capabilities and confidence of an LLM to its users (Spatharioti et al., 2023), or introducing various types of frictions that the user must bypass to receive the advice (Collins et al., 2024). Since interventions are being developed and evolving at a rapid pace, many of these methods lack comprehensive evaluations of appropriate reliance in real users, especially those focusing on technical contributions (Lin et al., 2022; Becker and Soatto, 2024; Bowman et al., 2022). Furthermore, people’s subjective perceptions—such as their self-confidence when receiving advice from LLMs—can have a complex relationship with their reliance behaviour (Kim et al., 2024b; Cao and Huang, 2022; Lu and Yin, 2021; Si et al., 2023), which warrants further exploration. Finally, the impact of different intervention types on reliance and perceptions remains unclear, as existing studies do not benchmark multiple interventions. We seek to address these gaps by asking the following research questions:
@@ 2.1. Appropriate Reliance on AI
Factors that influence reliance include the expertise of the user in the task, their cognitive engagement in the task, and the information presented by the AI (Chen et al., 2023; He et al., 2023; Gajos and Mamykina, 2022). On the individual level, people may also be averse or appreciative towards algorithms, perhaps due to pre-conceptions or judgments formed after seeing the AI’s performance (Kim et al., 2023; Dietvorst et al., 2015; Lu et al., 2023; Nourani et al., 2021; Ma et al., 2023). A number of strategies have been developed to augment users’ engagement and to provide them with more information of the AI decision, such as through providing explanations (Reingold et al., 2024; Vasconcelos et al., 2023b; Zhang et al., 2020; Schoeffer et al., 2024). The best interventions for appropriate reliance should be designed to support the users’ abilities to verify the AI’s recommendation, not necessarily to provide a fully transparent system (Fok and Weld, 2023; Poursabzi-Sangdeh et al., 2021; Wang and Yin, 2021). We adapt appropriate reliance metrics (Schemmer et al., 2023) to assess the ability of LLM interventions to help users verify outputs and increase appropriate reliance.
@@ 2.2. LLMs as Decision Aids
Recent research has investigated the role of LLMs in human-LLM collaborations for applications such as writing (Liang et al., 2024), communication (Fu et al., 2024), coding (Kazemitabaar et al., 2023; Nam et al., 2024), education (Lo, 2023; Kumar et al., 2023b; Bo et al., 2024), fact-checking (Wolfe and Mitra, 2024; Kim et al., 2024a; Si et al., 2023), information search (Spatharioti et al., 2023; Salikutluk et al., 2023; Sharma et al., 2024), and other specialized domains (Choudhury and Shamszare, 2023; Cheong et al., 2024; Yang et al., 2023). In these tasks, the LLM can provide guidance and information for the user to assist in their decision-making, learning, and creations. They can also offer antagonistic advice to induce the user into considering alternative options (Chiang et al., 2024; Ma et al., 2024a; Cai et al., 2024). Even when the user is just reading without explicitly searching for advice, the opinions embedded in the LLM outputs can be influence the user’s own opinions (Ferguson et al., 2023; Sharma et al., 2024; Krügel et al., 2023). In technical AI research, the quality of LLM responses has been improved through methods such as chain-of-thought with intermediate reasoning steps (Wei et al., 2022), debating amongst multiple LLMs (Khan et al., 2024; Kim et al., 2024a), and enforcing self-consistency in outputs (Wang et al., 2022; Becker and Soatto, 2024).
In ideal cases, the human-LLM team can reach better performance as a team than individually, tapping into the strength of their combined knowledge and skill sets (Bowman et al., 2022). However, the question of how to induce appropriate reliance is still open. While using an LLM for information search can be just as efficient as a search engine, the user’s ability to distinguish incorrect information is diminished, resulting in a higher rate of over-reliance on incorrect information (Si et al., 2023; Xu et al., 2023). On the other hand, aversion to LLM advice has also been observed, resulting in users losing performance gains that they could have achieved through relying on the LLM (Zhang, 2023).
@@ 3. Methodology
Two figures labelled a) diagram depicting the experimental procedure as explained in Methods, and b) sample of LSAT task question with the corresponding advice that each of the four conditions saw.
We designed a large-scale, randomized, and incentivized experiment to evaluate the efficacy of three reliance interventions applied to the outputs of GPT-4o111We used gpt-4o-2024-08-06 as it was the most state-of-art GPT model at the time of experimentation. in two question-answering tasks. To answer our research questions, we compare the effect of the interventions on performance, reliance behaviour, confidence calibration, and time of engagement against the effect of the baseline Control condition. In this section, we outline our choices of tasks, procedure, measures, and conditions.
@@ 3.1. Question-Answering Tasks
LSAT Logical Reasoning. Based on preliminary results from piloting multiple-choice questions from several domains, we found that logical reasoning questions best balanced difficulty with accessibility. We selected LSAT logical reasoning questions222Law School Admission Test® (LSAT) questions created by Law School Admission Council (LSAC). as the domain for the multiple-choice task and adapted two multiple-choice LSAT logical reasoning questions, both labeled ‘difficult’. Each LSAT question has five options, of which exactly one is correct (determined based on LSAT answer keys). We generated both correct and incorrect LLM advice for each question for use during the experiment. A sample question and corresponding LLM advice for all four conditions can be found in Figure 2(b), and screenshots of what participants saw can be found in Figures 1(b) in Appendix A.
@@
While LSAT questions can be difficult to comprehend, they are designed to test for aptitude in law school rather than knowledge of the legal domain. As such, they do not contain undefined jargon, nor do they require domain knowledge (satisfying D1). In addition to reasoning, LSAT questions also test analytical and reading comprehension skills, making them non-trivial to answer (satisfying D2). Furthermore, we reworded a question so that entering it into a search engine would not result in the correct answer (satisfying D3). Finally, they lend our study further external validity by emulating situations where people use LLMs to assist them with text comprehension or reasoning, in addition to people studying for tests such as the LSAT. While the MMLU dataset is another popular multiple-choice question benchmark for LLMs (Hendrycks et al., 2020), we did not find it suitable since it is skewed towards fact-based, know-it-or-you-don’t questions.
@@
Numerical Estimation. While most previous studies in human-AI collaboration focus on multiple-choice or binary outcome questions, we believe that introducing a Numerical Estimation (NE) task adds a valuable new dimension to this research space. Among the various forms of NE tasks, such as predicting values in regression tasks, we opted for an image-based format based on a classic carnival game: estimating the quantity of beans in a glass jar. We posit that this is a gamified task that is familiar to most people and does not require advanced mathematics (satisfying D1). Despite the simplicity of the estimation process, it is still challenging to achieve accurate results (satisfying D2). Lastly, since we took the pictures and many of the beans are occluded, it is impossible to use external assistance to “cheat” on this task (satisfying D3).
@@
For the experiment, we captured photos of glass jars with manually-counted quantities of beans from both side perspectives and bottom perspectives. One set featured a smaller jar containing 403 chickpeas, while the other contained a larger jar with 856 jellybeans. The pictures and sample LLM advice can be found in Figures A.2-A.3 in Appendix A. Unlike in multiple-choice tasks, where advice can be objectively classified as good or bad, the goodness of the advice in the Numerical Estimation setting depends on the participant’s initial estimate. For example, while advice recommending a number that is far lower than the ground truth is objectively wrong, it is still better than an even lower participant estimate. To account for this variability, we provided five distinct levels of LLM advice relative to the ground truth quantity: very low, low, good, high, and very high. We chose these levels of advice based on results from initial pilot experiments to balance the proportion of bad advice with good advice. Visualizations of the advice levels with the distribution of participant guesses are shown in Figure C.1 in Appendix C.
@@ 3.2. Procedure
In the experiment, each participant was randomly assigned to one of four conditions and then answered two questions from the LSAT task and two questions from the Numerical Estimation task, randomized so that the LSAT phase and Numerical Estimation phase were equally likely to appear first. The orders of questions within each category were also randomized. Participants received pre-generated advice from GPT-4o. This advice was randomly sampled for each question, meaning participants could receive any mix of ‘good’ or ‘bad’ advice. After completing the tasks, participants completed a short post-survey. See Figure 2(a) for the full diagram of the experimental flow and Figure 2(b) for sample advice for all four conditions for an LSAT question (for the advice provided in an Numerical Estimation question, see Figure B.2).
@@
Incentivization: Real-world decision-making involves balancing risk and reward. To simulate these stakes and induce incentive-compatible behaviour, we introduced a bonus, amounting to a maximum of 20% of the total participation pay. Participants earn rewards for correct answers, with two opportunities for bonus pay: one before, and one after receiving LLM-generated advice. In the LSAT questions, participants received a bonus if they selected the correct answer, while for the Numerical Estimation questions, participants earned a bonus if they guessed within 10% of the true count.
@@ 3.3. Measures
Flowchart demonstrating reliance pathways. There are three categories of elements: Initial Outcome, Action, and Reliance Behaviour. The diagram depicts the categories of Negative Self-Reliance (NS), Positive LLM Reliance (PL), Positive Self-Reliance (PS), and Negative LLM Reliance (NL) as described in the paper. Regardless of Action, “LLM advice is the exact same as Initial Answer” flows to “No reliance”.
What does it mean to rely on LLM advice? In multiple-choice questions, we follow previous work by defining reliance as changing one’s answer to the LLM’s suggestion. In numerical estimation questions, we define reliance as an answer with positive weight-of-advice (WoA), where WoA is calculated as — in short, this means that the participant was influenced by the LLM’s advice to adjust their estimate. We further define the reliance outcomes as falling into one of four categories:
@@
Negative LLM Reliance (NL): the participant changes their original answer to follow the LLM’s advice, but worsens or fails to improve their performance. This is also referred to as over-reliance.
Negative Self-Reliance (NS): the participant does not follow the LLM’s advice, but doing so would have improved their performance. This is also referred to as under-reliance.
PL and PS can be grouped together as Correct Reliance Behaviours, while NL and NS are Wrong Reliance Behaviours. NL represents over-reliance on LLM advice (the participant took the advice when it worsens or fails to improve improve their performance) and NS represents under-reliance (the participant didn’t take the advice but should have). We note that PL is possible even if the LLM advice is incorrect, as long as there is a notion of relative error—such as switching to a less wrong answer suggested by the LLM, which applies to the Numerical Estimation task and can extended to other real-world tasks. In contrast, the LSAT task lacks relative scoring across options, treating all incorrect answers as equally wrong; so relying on incorrect advice counts as over-reliance. This disambiguation is made since the LLM advice would not improve performance, so it would be preferable for the user to engage in the task and self-rely, allowing them to learn from the experience (Gajos and Mamykina, 2022). From these behaviour classifications, we compute the following appropriate reliance metrics:
Relative LLM Reliance (RLR) = Rate of correctly relying on the LLM when the LLM advice is helpful.
Relative Self-Reliance (RSR) = Rate of correctly relying on oneself when the LLM advice is harmful.
Appropriate Reliance Ratio (ARR) = Ratio of correct reliance on helpful advice to wrong reliance on harmful advice, where:
ARR ¡ 1: Higher rate of over-reliance than positive reliance, signifying inappropriate reliance behaviour.
@@
ARR ¿ 1: Higher rate of positive reliance than over-reliance, signifying appropriate reliance behaviour.
@@
Confidence and Subjective Perceptions. For each question, we collect the participant’s confidence in their answer prior to receiving LLM advice (Initial Confidence), and their confidence in their answer after receiving the LLM advice (Final Confidence) . Confidence is measured using a unipolar Likert scale from 1 (highly uncertain) to 5 (highly certain), which allows us to calculate the Change in Confidence as a result of receiving LLM advice. Ideally, well-calibrated confidence should result in lower confidence scores when participants make wrong reliance decisions. We compare the change in confidence between correct and wrong reliance outcomes for each of the conditions, where statistical differences are tested with Mann-Whitney U-test for non-normal distributions with significance threshold .
@@
ChatGPT’s advice made me more confident333Note that this measure of post-task confidence is different from the question-level confidence (the target of RQ2) described previously..
@@
Task Performance. Since our experiment is not designed to emulate the realistic performance of LLMs in tasks, we do not particularly emphasize the human-LLM collaborative performance. However, we explore the initial error from participants independently answering questions, and compare this with their updated error after re-answering questions with LLM advice. We also compute a Best Possible Performance metric, which simulates the outcomes as if the human made the correct reliance decision in every single question, representing the best possible accuracy of a human-LLM team.
Time of Engagement. To understand participant engagement with the LLM’s advice, we measure the amount of time that participants spend on each stage of question-answering. From these two time measurements, we calculate the relative time as the ratio between time taken to engage with the advice, to the time taken to answer the initial question (Relative Time). A larger value potentially signifies higher engagement in the advice, which may correlate with the complexity of the intervention paired with the advice.